y=-x-5 fin the slope and graphing
Answer by mananth(16946)

y=-x-5
compare with y= mx+b
m=-1 slope = -1
find the x&y intercept points
x=0,y=-5 (0,-5)
y=0,x=-5(-5,0)
Plot the points and draw the line graph